### v1

@@ 2 Electrostatic boundary conditions
Fig. 2 provides a schematic overview about the three main cases to setup electrostatic boundary conditions. To understand the impact of the specific boundary condition we consider its ability to correctly describe a key feature of an electrostatic interface namely the energetics and reaction rates of electrochemical interactions. As prototype example we consider here the case of an initially neutral atom or molecule at the interface (marked as initial in Fig. 2) that is positively charged (marked as final) after the reaction. Examples for this type of electrochemical reaction can be the dissolution of a metal surface in contact with water (wet corrosion) marcus2009 ; Huang2020Mg ; Yuwono2019 or a reaction step in an electro-catalytic reaction rossmeisl ; janik2009 ; laasonen ; santos2022 . The driving forces for this reaction are on one side the large solvation energy an atom or molecule can gain when becoming charged and/or the presence of the strong electric fields at such interfaces which energetically favor the movement of charge along the field.
Figure 2: Schematic representation of the possible electrostatic boundary conditions, as well as their associated electric ($\rho_{E}$) and displacement ($\rho_{D}$) field distributions. The circles with positive (+) and negative (-) charges depict a prototypical reaction. The green shaded area in the electric field distribution plot shows the region within which the reaction can occur.
@@
A product of the here considered prototype reaction, namely a single atom which is (singly) positively charged after the reaction, is an extra electron, which initially will stay on the surface. A key question when modeling such a reaction is what happens to this extra charge on the surface and are the boundary conditions adequate to reproduce the behaviour of the macroscopic system?
As a first scenario we consider constant charge conditions. In this case, the net charge on the electrode where the reaction takes place and on the counter electrode (not shown here) is kept fixed and its sum is zero, i.e., the total setup of electrode and counter electrode is charge neutral. This scenario corresponds to constant electric displacement $D$ and is sketched in Fig. 2a. At constant charge conditions the number of excess charges (shown as white circles with a plus at the electrode) is fixed. Consequently, the extra electron created by the charge transfer reaction cannot be compensated and forms together with the positive ion an electric dipole along the surface normal. For a supercell geometry with periodic boundary conditions, this dipole is infinitely repeated laterally giving rise to an effective capacitance and thus a large surface dipole. As a consequence, the average potential $\Phi\right>$ will be significantly shifted away from the initial voltage.
[…]
@@
We note, however, that for computing equilibrium quantities - such as the potential of zero charge (PZC) - for a surface in contact with an electrolyte, the constant charge approach ($D=$ const) for the special condition of zero charge is well suited and has been successfully employed in several studies surendralal2021 ; pzc_jle . The reason why these boundary conditions work well in this case is that in the absence of chemical reactions the surface structure remains unchanged and the PZC is characterized by the condition that no net charge is transferred to or away from the electrode, which is automatically given by the condition $q=$ const.
@@
The second scenario assumes that the voltage $\Phi$ is kept constant, i.e., fluctuations in the potential are not possible ($\Phi\equiv\left<\Phi\right>$), see density distribution for $E$ in Fig. 2b. To keep the potential constant, charges can be exchanged (i.e. flow back and forward) with the potentiostat. Thus, for the model reaction considered here, once the extra electron is at the electrode a charge $q$ will be transferred until the shift between initial and final potential becomes zero. Therefore, these boundary conditions provide a realistic description of the electrochemical current that flows when keeping the voltage constant and charge transfer reactions take place. Experimentally, this would be measured as current-voltage (U-I) characteristic. While the resulting net current can be described, the lack of the thermally induced fluctuations in the electric field implies a qualitatively different reactive area. As an analogy, in the theory of thermostats it is well understood that “[…] for small systems or when the observables of interest are dependent on the fluctuations rather than on the averages […]csvr , such a method cannot be used. For polar systems such as water, thermal fluctuations cause fluctuations in the net interface dipole, which lead to fluctuations in the potential $\Phi$. Since these voltage fluctuations represent the actual physics of the system, it is important to not constrain them artificially by the condition of constant voltage (cf. density distribution of $E$ in Fig. 2b). Due to the absence of fluctuations, electric fields that are more beneficial for the reaction are absent.
Since the voltage distribution for typical DFT supercell sizes is in the order of a few eV, whereas the characteristic thermal energy at room temperature $k_{B}T$ is only 25 meV, Boltzmann population will occupy mainly the energetically low-lying states, which are missing at constant voltage (cf. Fig. 2b).
The most general boundary conditions for the supercell are the thermodynamically open ones, both with respect to charge transfer and fluctuations, as well as to voltage control (e.g. $\Phi\right>=$ const.), cf. Fig. 2c. These boundary conditions are able to describe the electrochemical net current in analogy to the constant voltage approach, but without any artificial constraints on either the field or the charge fluctuations. Unlike the $D=$ const and $E=$ const approaches, where the reactive area is either changing or entirely missing, these open boundary conditions preserve the reactive area before and after the reaction (see distributions in Fig. 2c).
@@
Fully open boundary conditions imply, however, that the system is able to perform external work $\Delta W=\Delta q\cdot\Delta E$, where $\Delta q$ and $\Delta E$ are the change in the charge and the electric field, respectively. In the first two scenarios (cf. distributions in Figs. 2a,b) either $\Delta q$ or $\Delta E$ is artificially enforced to be exactly zero, so that $\Delta W\equiv 0$. Under open boundary conditions, however, $\Delta W$ becomes finite. The finite $\Delta W$ permits the system on short time scales to “borrow” energy from the ensuing charge and potential fluctuations, in analogy to the theory of thermostats csvr . This ability to perform external work is necessary in order to preserve the reactive area before and after the reaction, as illustrated in Fig. 2c. On the other hand, the finite $\Delta W$ implies that introducing potential control via relaxing $\Phi(t)$ towards its targeted mean $\langle\Phi\rangle$ dissipates energy (represented schematically by the resistor in Fig. 2c).
@@
These fully open boundary conditions otani2012PRL ; deissenbeck2021 accurately reproduce in a supercell sized box the behavior of voltage and charge measured within the full macroscopic system. These conditions are thus able to describe all aspects of an electrochemical system including electrochemical reactions or dynamic changes in the surface or adatom structure. Only when computing the intrinsic electrode potential of a system, constant (zero) charge boundary conditions may be used.
3 Design of a computational counter electrode
Realizing potential control - as outlined above - in electronic structure simulations, requires the ability to apply an electric field or introduce a charge on the electrode surface, so that the field distribution inside the supercell accurately describes the real physical system (cf. atomistic scale in Fig. 1). To avoid a singularity in the electrostatic energy when using supercells with periodic boundary conditions the cell must be charge neutral. As a consequence, the excess charge at the surface must be exactly compensated by an opposite charge. If no explicit compensating charge is added, DFT codes drop the $G=0$ plane wave component in Fourier space, resulting effectively in a spatially homogeneous background charge which exactly compensates for the excess charge at the surface.
Figure 3: Comparison between the potential in the macroscopic system and the potential obtained in DFT supercell simulations via a computational electrode. The potentials agree within the microscopic electrochemically active region. The fraction of the double layer charge that is located beyond the boundaries of the supercell must be placed on the DFT computational electrode to ensure charge neutrality. Therefore, at larger distances both potentials differ by a constant amount. A suitable correction can be performed a posteriori.
To provide more control in achieving realistic field distributions, a computational counter electrode is included. We emphasize that the computational counter electrode does not play the role of any physical electrode (i.e. neither the physical counter electrode nor the reference electrode). Its sole purpose is to compensate for the charge at the electrified surface, to introduce the desired field distribution in the microscopic region targeted by the simulations, and to provide the boundary conditions set by the electrochemical environment that is external to the simulation cell. In other words, the computational counter electrode accounts for the residual charge of that part of the electric double layer, that is located outside the simulation cell. Therefore, inside the simulation cell, the field and potential distributions of the simulated system are equal to the ones of the macroscopic physical system within the electrochemically active region. Outside the simulation cell, the potentials differ by an offset, cf. Fig. 3. Suitable corrections can be straightforwardly derived and applied a posteriori via purely electrostatic modelling without the need for simulations at the electronic structure level of theory.
Realizing a suitable computational counter electrode is a challenging task. Most density-functional theory (DFT) simulations for electrode-electrolyte interfaces are performed using supercell calculations in periodic boundary conditions (PBCs) che1 ; filhol2006 ; taylor2006 ; schnur ; gaigeot ; Galli2021 ; cyli . The use of PBCs, which is required in order to describe extended interfaces, strictly enforces (i) simulation cells that are charge neutral in total (due to Gauss’s law in electrostatics) and (ii) electric field distributions that obey the periodicity of the supercell. Moreover, (iii) in available DFT codes there is only a single Fermi level that is constant throughout the supercell. It is thus explicitly impossible to maintain a working electrode and a counter electrode within the same supercell at different Fermi levels in standard DFT simulations.
Therefore, while it is trivial in the case of thermostats or barostats to adjust the control parameter (i.e. the kinetic energy or the cell volume, respectively), this is considerably more difficult in the case of potentiostats (electrode charge). Multiple approaches to apply fields or charges in supercells have thus been explored over the past decades, to address the restrictions outlined above. Fig. 4 provides a schematic overview of the main computational setups.
One of the earliest approaches was suggested by Neugebauer and Scheffler neugebauer1993 , extending the dipole correction neugebauer1992adsorbate to place a slab in a capacitance. Later, Lozovoi and Alavi lozovoi used a symmetric slab setup, cf. Fig. 4a, where the surfaces are charged by introducing a net charge to the system. In order to ensure charge neutrality of the super cell, an equal and opposite homogeneous background charge is used as a computational counter electrode. While this setup can be easily applied in any DFT code, it is unable to provide a direct measure for the electrode potential: in principle, the electrode potential is given by the derivative $\Phi=\frac{\partial\mathcal{H}}{\partial q}$ of the DFT total energy $\mathcal{H}$ with respect to the charge $q$, which is, however, a notoriously difficult quantity to compute in 3D PBC supercell calculations defect_review ; FNV_PRL . Also, the distribution of the counter-charge as homogeneous background is very different from the distribution of the counter-charges in a realistic electrochemical cell. In the latter, the counter-charges are realized by the counter-electrode and the net-charge of the solvated ions in the electrolyte, which show a distribution very different from a spatially homogeneous and constant background.
Otani et al. therefore proposed to use 2D PBCs to describe the lateral periodicity of the interface, while lifting the PBC parallel to the surface normal via Green’s function techniques otani2006esm ; greens ; otani2013esm2 , cf. Fig. 4b. Alternatively, one may lift the PBCs selectively only for the electric field in the supercell via electric enthalpy-based approaches pasquarello2002 or Berry phase techniques resta and the modern theory of polarization stengel ; stengel2 , cf. Fig. 4c. Here, the electrostatic potential drop across the supercell is described by a geometric phase, which accumulates when moving from one periodic image of the supercell to the next. Such approaches to describe electric fields in supercells are demanding to implement. Hence, they are available in only some of the $\sim$ 40+ publicly available DFT codes sandia . Moreover, the practical implementations of this approach work typically only for systems with a bandgap. Thus, only systems containing semiconductor or insulator electrodes can be studied.
Figure 4: The various approaches used to achieve a charged (electrified) electrode-electrolyte interface within a DFT supercell. Electrodes are shown as grey rectangles at the edge(s) of the supercell (marked by a dotted line) and the electrolyte as a blue area within it. The red hue covering the supercell in a) represents the homogeneous background charge. Ions and charges at the electrodes are shown as spheres containing a plus or minus, while atoms are shown as coloured spheres. The resulting electrostatic potential curve is depicted by the red line.
For this reason, minimally invasive approaches working within the constraints of 3D PBC have been exploited, as well. Multiple groups proposed gche1 ; gche2 ; agross2020 ; omranpoor2023 to vary the number of H+ ions contained within the DFT supercell grand-canonically. For the same purpose, Cucinotta et al. cucinotta introduced an intentionally unbalanced number of cations and anions in the electrolyte. Both approaches give rise to an uncompensated net-charge. Since the total system is charge neutral, the non-compensated charge accumulates at the interface, cf. Fig. 4d. By construction, this charge can assume only integer values. Even when using rather large supercells, the number of possible charge states at the interface is rather limited. This limited number of charge states at the interface severely restricts the variability of the ensuing electric fields and, hence, the applicability of techniques to sample thermodynamically open boundary conditions with $\langle\Phi\rangle=\text{const}$.
A continuous variation of the charge on the working electrode is achieved via charged implicit solvents mathew2014 ; mathew2019 . Also, hybrid solvation schemes have been suggested, where a region of explicit electrolyte is complemented by a charged implicit solvent or continuum dielectric goldsmith ; Melander2024 , which are placed at the edge of the explicit electrolyte region. Hybrid approaches of this type require an implementation within the respective electronic structure package. For this reason, Surendralal et al. surendralal2018 proposed to replace the charged implicit region by an explicit atomistic counter electrode. This approach uses pseudo-atoms, i.e. neutral objects with fractional proton numbers shiraishi , to charge the working electrode. The employed setup introduces an explicit counter electrode consisting of Ne pseudo-atoms and a vacuum region into the supercell, and is shown in Fig. 4e. Apart from enabling a continuous variation of the charge within the supercell, this approach has also the advantage, that: (i) it avoids the use of a compensating background, since the overall supercell remains neutral, (ii) all its design elements can be easily realized in a standard DFT code and do not require changes to the code and (iii) it works equally well with metallic and semiconducting electrodes. Moreover, Surendralal et al.’s setup is straightforwardly extended to describe fully open boundary conditions for the electrode charge deissenbeck2023 . Dudzinski et al. dudzinski2023 recently extended the idea of using fractionally charged pseudo-atoms to modify the total charge of the water molecules in an aqueous electrolyte, although at the price of relinquishing control over the distribution of charges within the electrolyte.
Charging the supercell    & Application to solid/liquid interfaces
Method Charging Geometry Boundary Electrode
slab solv conditions potential
[…]
Sheet charge/jellium lozovoi discrete sym vac const $D$ yes
[…]
       filhol2015 impl yes
Electrolyte with net charge
[…]
Table 1: Methods to charge supercells and model electrochemical solid/liquid interfaces, and their key properties/approximations.
We emphasize that modelling the computational counter electrode either as a charged continuum region outside the explicit electrolyte or as an object consisting of actual atoms creates a hydrophobic gap between the computational counter electrode and the electrolyte, leading to an undesired and non-negligible potential drop within this region deissenbeck2023 . Moreover, this interfacial potential loss at the computational counter electrode increases the likelihood of reaching a dielectric breakthrough yoo2021 as the field strength increases (see also discussion in section 4).
To eliminate the hydrophobic gap at the computational counter electrode and any associated potential loss, in a newer development deissenbeck2023fse the setup proposed by Surendralal et al. surendralal2018 has been extended to the use of a fully solvated electrode, cf. Fig. 4f. In this approach a Gaussian counter charge, placed within the aqueous region, is used to charge the working electrode. While this method requires slight modifications in the DFT code, it enables considerably larger potential drops and realistically strong electric fields at the working electrode/water interface. Moreover, future developments will enable an accurate description of the diffuse region at the interface by exchanging the Gaussian charge for a selfconsistently determined counter charge distribution, that accurately describes the charge density distribution of the double layer in that region.

### v2

@@ 3 Electrostatic boundary conditions
In the preceding section, we gave an overview about available surrogate models which replicate the impact of the electrochemical double layer and the real system’s potential distribution inside the supercell. A critical aspect of such a construct is selecting appropriate electrostatic boundary conditions for the system, which have to include the working electrode, explicit electrolyte, and surrogate model. It is important to note that different regions of the electrochemical cell are governed by distinctly different time scales. The double layer relaxes slowly on nanosecond timescales due to the relatively slow ionic motion santos2024 , while the electrode surface equilibrates rapidly, on timescales governed by electron mobility melander_dyn . The latter timescale is orders of magnitude faster. The chosen electrostatic boundary conditions must describe both the fast equilibration of the electrode charge and the slow relaxation of the double layer (via the surrogate model) equally well. This is particularly relevant under reactive conditions, where charge transfer processes must be captured, including the energetics and reaction rates of electrochemical processes.
@@
As a prototype example, we consider here an initially neutral atom or molecule at the interface (cf. Fig. 4a) that becomes positively charged after a reaction. Such processes are a crucial step in metal dissolution in contact with water (e.g., wet corrosion) marcus2009 ; Huang2020Mg ; Yuwono2019 or in an electrocatalytic reaction rossmeisl ; janik2009 ; laasonen ; santos2022 . The driving forces for these reactions are the large solvation energy gains upon charging the solute and the strong electric fields at the interface that favor charge movement along the field. A product of the here considered prototype reaction, namely a single atom which is (singly) positively charged after the reaction, is an extra electron, which initially will stay on the surface. A key question when modeling such a reaction is what happens to this extra charge on the surface and are the boundary conditions adequate to reproduce the behaviour of the macroscopic system?
To explore the validity and performance of different boundary conditions, we performed MD simulations for a typical DFT sized supercell with lateral dimensions of 10 Å by 10 Å, containing 64 water molecules. We measured the voltage drop between the working electrode and regions close to the electrode surface ($U_{g}$), and deeper inside the electrolyte ($U_{i}$) flucts , respectively, mimicking the placement of reference electrodes, cf. Fig. 4a. The resulting potential distributions provide a direct insight into the short-range electrostatics near the electrode and the long-range behavior in the electrolyte, and allow thereby testing the validity of the commonly discussed and applied boundary conditions under reactive conditions.
As a first scenario we consider constant charge conditions. In this case, the net charge on the electrode, where the reaction takes place, and the counter charge are kept fixed, and their sum is zero, i.e., the total setup of electrode and surrogate model is charge neutral. This scenario corresponds to constant electric displacement $D$ at the electrode-electrolyte interface. In the absence of charge transfer, the value of $D$ is directly related to the surface charge density $\sigma$ at the interface with $D=\sigma$.
Figure 4: a) Schematic representation of the supercell with an applied voltage between the working electrode and the surrogate double layer model. Voltages are measured between the working electrode and regions in close vicinity to the electrode surface ($U_{g}$) or deeper within the electrolyte region ($U_{i}$), respectively. b/c) Measured voltage distributions for different electrostatic boundary conditions flucts , see text. The height of the delta peak in b) is rescaled for better visibility.
Under constant charge conditions (constant displacement), the number of excess charges on the electrode surface (represented as a white circle with a plus) remains fixed. This constraint is reflected in the distribution of the voltage $U_{g}$ between the working electrode and the region directly above its surface. In the absence of charge transfer to or from the working electrode, the total surface charge is conserved. Consequently, the voltage drop across this region remains constant, leading to a delta distribution for $U_{g}$ (blue line, Fig. 4b).
Once the charge transfer reaction proceeds, the resulting extra electron remains localized on the electrode surface and cannot be compensated for under constant charge conditions. This unbalanced electron, in conjunction with the positively charged ion in the electrolyte, forms an electric dipole oriented along the surface normal. For a supercell geometry with periodic boundary conditions, this dipole is infinitely repeated laterally, giving rise to an effective capacitance and thus a large surface dipole. Consequently, the mean of the voltage distribution $U_{g}$ shifts substantially from its initial value. This shift disrupts the electrostatic conditions necessary for further reactions of the same type, effectively halting the process.
This behavior illustrates a limitation of constant charge boundary conditions when simulating reactive processes, where the interface dipole can change, e.g., due to electrochemical reactions, changes in the surface or adsorbate geometry, etc. In a real system, the electrode surface charge equilibrates rapidly on timescales governed by the electron mobility. Thus, the real system does not feature a delta distribution for $U_{g}$, but obeys a Gaussian (normal) distribution with a deviation of $k_{B}T/C$ deissenbeck2021 ; deissenbeck2023 , where $C$ is the effective (average) capacitance of the interface at the given mean potential. This capacitance characterizes the ability of the electrode to accommodate charge fluctuations and depends on both the interfacial geometry and the electronic structure. Therefore, approximating the surface charge as constant requires a careful analysis of the validity of this approximation for the specific study.
@@
We note, however, that for computing equilibrium quantities of a surface in contact with an electrolyte, i.e. where changes in the interfacial dipole are small – such as the potential of zero charge (PZC) or the interfacial capacitance – the constant charge approach ($D=$ const) for the special condition of zero charge is well suited and has been successfully applied in several studies surendralal2021 ; pzc_jle ; sundararaman2022 . The reason why these boundary conditions work well in this case is that in the absence of chemical reactions the surface structure remains unchanged and the PZC is characterized by the condition that no net charge is transferred to or away from the electrode, which is automatically satisfied by the condition $q=$ const.
Moreover, one should keep in mind that the double layer relaxes relatively slowly, on the order of up to nanoseconds, due to the sluggish motion of ions in the electrolyte santos2024 . While constant charge boundary conditions fail to accurately describe the dynamics of the voltage $U_{g}$ close to the electrode surface, they are well-suited for capturing the dynamics of the voltage $U_{i}$ within the double layer region of the cell through the surrogate model. As shown in Fig. 4c (blue curve), the thermal motion of the electrolyte induces polarization fluctuations, which are correctly reproduced under constant charge boundary conditions.
@@
In order to lift the inability of constant displacement (charge) schemes to account for the charge dynamics of the working electrode under reaction conditions, a second approach suggests to keep the voltage between the working electrode and the surrogate model precisely constant, i.e., fluctuations in the applied voltage are not possible ($U\equiv\left<U\right$>). To keep the potential constant, charges can be exchanged (i.e. flow back and forward) with the potentiostat. Thus, for the model reaction considered here, once the extra electron is at the electrode, a charge $q$ will be transferred until the shift between initial and final voltage across the cell is eliminated. Therefore, these boundary conditions provide a realistic description of the electrochemical current that flows when keeping the voltage constant and charge transfer reactions take place. Experimentally, this would be measured as current-voltage (U-I) characteristic.
Constant electric field boundary conditions are particularly effective in capturing the fast charge dynamics of the working electrode surface. Unlike the delta distribution obtained for $U_{g}$ under constant charge conditions (Fig. 4b, blue curve), these boundary conditions yield the expected Gaussian distribution for $U_{g}$ (Fig. 4b, red curve), accurately reflecting thermal fluctuations near the electrode. However, constant electric field conditions enforce an instantaneous relaxation of the double layer, as the voltage between the working electrode and the surrogate model is held precisely constant at all times. This assumption deviates significantly from the behavior of the real system, where ionic relaxation occurs on much slower time scales. As a result, the voltage distribution $U_{i}$ within the electrolyte narrows considerably when approaching the double layer region (Fig. 4c, red curve). For DFT-sized supercells, this distribution is much narrower than the one obtained under constant charge conditions, which approximates the double layer as a static entity.
As expected, constant charge boundary conditions are better suited for describing the slow relaxation dynamics of the double layer, while constant field boundary conditions more effectively capture the rapid charge dynamics of the electrode surface. However, neither approach is capable of accurately representing both aspects simultaneously, failing to reconcile the short-range behavior near the electrode with the long-range behavior deeper inside the electrolyte.
We emphasize that the choice of boundary conditions can significantly influence the predicted reaction mechanisms and rates in electrochemical simulations. For typical supercell sizes used to model the electrochemical interface, the average voltage across the electrode electrolyte interface can fluctuate by approximately $\Delta\Phi\sim\pm 1$ V PhD_sudarsan ; rossmeisl ; spohr , as illustrated in Figs. 4b/c. These fluctuations directly impact the reaction barrier and, consequently, the reaction rate.
The reaction rate depends on three distinct timescales: (i) the timescale of the reaction itself ($\tau_{\text{reaction}}$), (ii) the timescale of the electrostatic potential fluctuations ($\tau_{\Phi}$) due to the fast equilibration of the electrode charge, and (iii) the much slower timescale over which the double layer equilibrates ($\tau_{\text{eq}}$). The latter includes solvent and ionic reorganization, such as dielectric relaxation of water and ion migration, and sets the rate at which the electrostatic environment around the reacting species can adjust. If $\tau_{\Phi}\gg\tau_{\text{reaction}}$, the reaction occurs under a quasi-static potential, and the rate reflects the instantaneous barrier. Conversely, if $\tau_{\Phi}\ll\tau_{\text{reaction}}$, the system averages over the potential fluctuations during the reaction. When $\tau_{\Phi}\sim\tau_{\text{reaction}}$, the dynamics of the fluctuations must be explicitly considered, as the reaction rate becomes sensitive to the correlation between potential fluctuations and barrier crossing.
In general, the reaction rate $w_{\text{reaction}}$ can be expressed as
$w_{\text{reaction}}\propto\int_{-\infty}^{\infty}P\left(\Phi\semicolon\;\tau_{\Phi},\tau_{\text{reaction}},\tau_{\text{eq}}\right)\cdot e^{-\frac{\mathcal{E}_{\text{barrier}}\left[\Phi\right]}{k_{B}T}}d\Phi,$
where $\mathcal{E}_{\text{barrier}}$ is the reaction barrier as a function of the instantaneous potential $\Phi$, and $P$ is the probability density of the system to adopt a specific electrostatic potential $\Phi$. This expression isolates the potential-dependent contribution to the reaction rate for simplicity. A full treatment would also require integration over molecular positions and momenta, which are omitted here. We emphasize that this timescale discussion is conceptual; accurate rate calculations require advanced sampling techniques such as umbrella sampling umbrella , metadynamics metadyn1 ; metadyn2 , forward flux Allen2006 ; Allen2009 , or the blue moon ensemble bluemoon . Importantly, the probability density $P$ – and hence the predicted reaction kinetics – depends on the chosen boundary conditions, as these influence both the magnitude of potential fluctuations and the timescales over which they evolve.
The most general boundary conditions for simulating electrochemical supercells are thermodynamically open, allowing for both charge transfer and dynamical fluctuations while maintaining voltage control. In this framework, the applied voltage is fixed only on average (e.g. $\left<\Phi\right$>= const.), enabling fluctuations around the mean value. On short time scales, the electrode surface charge and the voltage drop across the interface can vary significantly, reflecting the interplay between the rapid dynamics of electrode charge equilibration and the slower relaxation processes of the electrochemical double layer. These boundary conditions aim to accurately capture the inherently dynamic nature of the electrochemical interface.
@@
Fully open boundary conditions imply, however, that the system is able to perform external work $\Delta W=\Delta q\cdot\Delta\Phi$, where $\Delta q$ and $\Delta\Phi$ are the change in the charge and the potential, respectively. Under constant displacement or constant field boundary conditions, either $\Delta q$ or $\Delta\Phi$ is artificially enforced to be exactly zero, respectively, so that $\Delta W\equiv 0$. Under open boundary conditions, however, $\Delta W$ becomes finite. The finite $\Delta W$ permits the system to “borrow” on short time scales energy from the ensuing charge and potential fluctuations. This mechanism is analogous to the energy exchange in thermostat theories, e.g. “[…] when the observables of interest are dependent on the fluctuations rather than on the averages […].” csvr . The ability to perform work is crucial for restoring the applied potential to its target value after events such as charge transfer or changes in surface structure, thereby maintaining reactive conditions within the simulation. On the other hand, the finite $\Delta W$ implies that introducing potential control via relaxing $\Phi(t)$ towards its targeted mean $\langle\Phi\rangle$ dissipates energy.
@@
These fully open boundary conditions otani2012PRL ; deissenbeck2021 accurately reproduce in a supercell sized box the dynamic behavior of voltage and charge measured within the full macroscopic system. These conditions are thus able to describe all aspects of an electrochemical system including electrochemical reactions or dynamic changes in the surface or adsorbate structure. Computing the voltage distributions near the electrode surface ($U_{g}$) and deeper within the electrolyte ($U_{i}$) confirms the effectiveness of thermodynamically open boundary conditions (Fig. 4a/b, green curves). At short range, they replicate the behavior of constant field boundary conditions, accurately describing the fast dynamics of the electrode charge. At long range, they converge to the behavior of constant displacement boundary conditions, correctly modelling the slower relaxation of the double layer flucts .
4 Open thermodynamic boundary conditions
The surrogate double layer models discussed in sections 2 and 3 enable AIMD simulations at finite applied electric field for supercells that are charge-neutral in total, but with a varying number of electrons $n_{e^{-}}$ per supercell and thus, by extension, a varying charge $n$ on the working electrode. Based on these concepts, we now turn to discuss the realization of open thermodynamic boundary conditions with respect to the electrode charge $n$, as outlined in section 3.
Any such control scheme relies on a conjugate pair of extensive and intensive quantities - such as entropy and temperature (thermostat), or volume and pressure (barostat) - where the extensive quantity is used to drive the fluctuating intensive one to a desired target value [^14] We note that in the case of thermostats, the entropy is difficult to control in simulations. There is, hence, a need to perform a Legendre transformation in order to control the temperature via the particle kinetic energies instead. For barostats, in contrast, the volume is straightforwardly adjusted. In that respect, the construction of a potentiostat is conceptually much closer to a barostat than to a thermostat.. In the context of potential control, it is straightforwardly shown that the electrode charge $n$ and the electrode potential $\Phi$ – as well as the displacement $D$ and the field $E$ landau – form such a pair of conjugate quantities deissenbeck2021 . Since the electrode is connected to an external reservoir of charge, constructing a potentiostat amounts to deriving an equation of motion for the extensive charge $n_{e^{\text{-}}}(t)$, i.e. the exchange of charge between the electrode at potential $\Phi(t)$ and the external reservoir held at $\Phi_{0}$.
This task is complicated by the fact that the potentiostat is intended to describe the exchange of charge with a reservoir not just in static calculations, but in the context of molecular dynamics at finite temperature. The thermal motion of the electrons and ions, as well as the exchange of charge between the electrode and the reservoir, create thermal fluctuations in the electric field and - by extension - in the potential of the local system $\Phi(t)$. In the physical system, the identical argument applies to the external environment surrounding the microscopic region targeted by our simulations. As a result, the DFT supercell effectively resides within a fluctuating electric field generated by the thermal motion of electrons and ions in its local external environment. This embedding reflects the realistic coupling between the simulated region and its surroundings, essential for accurately modeling electrochemical processes.
It is the role of the potentiostat to mimic the contact of the supercell with this external environment. If the reservoir of charge were to be held at an exactly constant potential without any thermal fluctuations (effectively corresponding to $T=0$ K), the supercell would dissipate its thermal potential fluctuations into the external reservoir, thereby violating the 2nd law of thermodynamics. Such an artificial net energy drain would constantly drive the simulation system out of equilibrium. In order to prevent such a spurious energy transfer, the electric field fluctuations within the supercell must be in thermodynamic equilibrium with the field fluctuations of the external environment at the boundary of the supercell. Therefore, the external reservoir of charge itself must have a finite temperature.
The first such approach was introduced in 2012 by Bonnet et al. otani2012PRL , who drove the exchange of charge between electrode and reservoir via a Nose-Hoover algorithm. The Nose-Hoover approach is known to perform well for reasonably ergodic degrees of freedom. The electrode charge, however, is only a single degree of freedom and thus requires Nose-Hoover chains at the cost of an increased number of parameters (Nose-masses) and associated extra tuning. Moreover, the absence of suitable surrogate models for the double layer in 2012 led Bonnet et al. to construct their potentiostat for systems carrying a net excess charge. Hence, the Nose-Hoover potentiostat requires the derivative of the total energy with respect to the net excess charge as an input quantity. This quantity is notoriously difficult to obtain in DFT simulations. Those requirements proved to be serious obstacles in implementing the Nose-Hoover potentiostat into existing DFT codes.
In the context of potential control, driving the system potential $\Phi(t)$ towards the external potential $\Phi_{0}$ is the dissipation. Formally, in an equivalent circuit approach, the dissipation can be represented as a resistance $R$ between the part of the working electrode located inside the supercell and its surrounding external environment. Physically, the resistance $R$ describes the effective surface resistance of the electrode to dampen the small charge fluctuations at the electrochemical interface that arise at the surface due to thermal atomic vibrations, (electro-) chemical reactions etc. This damping is consistent with the fluctuation–dissipation theorem and the emergence of Johnson–Nyquist noise henkel16 . We emphasize that this is not a circuit-level resistance between electrodes. Much like the friction coefficient in a Langevin thermostat governs energy exchange between a particle and a thermal bath, the resistance $R$ controls the rate at which the electrode surface exchanges charge and potential fluctuations with its electrostatic environment.
This dissipation must then be balanced by the energy gained from the thermal potential fluctuations of the local external environment at the boundary of the supercell. In the equivalent circuit approach, the fluctuations are then given by the Johnson-Nyquist noise generated by the resistance $R$. It was shown in Ref. deissenbeck2021 , that the fluctuation-dissipation theorem in this case takes the form
where $f$ is the rate of change of the electrode potential $\Phi$, $\tau_{\Phi}$ is a relaxation time constant (typically in the order of 100 fs deissenbeck2023 ), $dW_{t}$ is the time differential of the Wiener stochastic process and $C_{0}$ the capacitance of the bare electrodes in the absence of a dielectric, i.e. $C=\epsilon_{r}C_{0}$ deissenbeck2021 ; deissenbeck2023 .
Such a thermopotentiostat was recently derived by Deißenbeck et al. deissenbeck2021 ; deissenbeck2023 via Ito integration gardiner of the fluctuation-dissipation relation shown above. Thereby, the thermopotentiostat provides a direct expression for the electrode charge $n$ at each discrete MD time step, under open thermodynamic boundary conditions. In contrast to the $D=\text{const.}$ and $E=\text{const.}$ approaches, no artificial constraints on either the charge or potential dynamics are imposed  [^15] In the limit of setting the equivalent resistance $R$ (or by extension the relaxation time $\tau_{\Phi}$) to either zero or infinity, respectively, electrostatic boundary conditions of $\vec{E}=\text{const.}$ or $\vec{D}=\text{const.}$ are recovered.. Instead, the thermopotentiostat is mathematically guaranteed to sample the charge and potential distributions depicted in Fig. 4, which maintain their averages even under charge transfer reactions. In addition, it no longer requires the derivative of the total energy with respect to the charge as an input quantity. Instead, $\Phi(t)-\Phi_{0}$ is the only mandatory input quantity. In computational setups such as the ones shown in Figs. 3e, g, h this potential difference can be conveniently measured via the total dipole moment - given, e.g., by the dipole correction neugebauer1992adsorbate - along the normal direction of the electrode surface. Therefore, in conjunction with any of those computational setups the thermopotentiostat can be straightforwardly included in any electronic structure package.
From a practical standpoint, whether a potentiostat is necessary depends on both the system size and the nature of the process being studied. For typical DFT simulations of electrochemical interfaces – often involving 100–1000 atoms and lateral cell dimensions of 1–2 nm – finite-size effects remain significant. In such systems, local fluctuations in the electrostatic potential, which can influence activation barriers, reaction selectivity, or adsorption geometries, are suppressed because the simulation cell is too small to capture the full range of independent microstates present in a macroscopic environment. Just as thermostats and barostats correct for the finite size of a system in temperature- or pressure-controlled simulations, the potentiostat serves to remove finite-size artifacts in voltage-controlled conditions. For this reason, and because the computational overhead is negligible, we recommend potential control in any simulation involving charge transfer, applied bias, or voltage-sensitive processes. In contrast, for closed equilibrium systems where no charge exchange occurs and the goal is, e.g., structural or spectroscopic characterization, ensemble control may be safely omitted – though the implications of the chosen boundary conditions should always be assessed. As the scale and realism of simulations continue to grow, careful consideration of ensemble control will remain essential for making physically meaningful predictions.